# An Automated Multi-modal Evaluation Framework for Mobile Intelligent Assistants Based on Large Language Models and Multi-Agent Collaboration

### v1

@@ Abstract.
journalyear: 2025copyright: rightsretainedconference: Companion of the the 2025 ACM International Joint Conference on Pervasive and Ubiquitous Computing; October 12–16, 2025; Espoo, Finland.booktitle: Companion of the the 2025 ACM International Joint Conference on Pervasive and Ubiquitous Computing (UbiComp Companion ’25), October 12–16, 2025, Espoo, Finlandisbn: 979-8-4007-1477-1/25/10doi: 10.1145/3714394.3756184ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Natural language processingccs: Computing methodologies Multi-agent systems
@@ 1. INTRODUCTION
  • Cross-platform validation across five major smartphone brands (vivo, OPPO, Xiaomi, Huawei, Honor), demonstrating practical applicability and universality;
@@ 4.2. Cross-Platform Evaluation Setup
Five major smartphone AI assistant platforms were selected representing diverse technological approaches: vivo’s BlueLM, OPPO’s Breeno, Xiaomi’s Xiao Ai, Huawei’s Assistant, and Honor’s YOYO, as well as three internet-based large model mobile apps (Doubao, Tencent Yuanbao, and DeepSeek Chat). To ensure fair comparison, input delivery protocols were standardized across all platforms, presenting each interaction scenario within the same time window using identical user accounts and device configurations.
@@ 4.4. Results and Analysis
First-order dimension Total Doubao Yuanbao deepseek Huawei Honor Xiaomi OPPO vivo
@@
Table 2. Each assistant shows obvious differences in different scenarios
@@
From the perspective of task type dimension (Table 2), each assistant shows obvious differences in different scenarios. In educational learning tasks, the overall performance of each model is better, with average scores generally exceeding 85 points. Internet assistants (such as Doubao 94.82 and DeepSeek 93.51) show strong multi-modal understanding in image-text problems. Mobile phone assistants demonstrate actual control ability, significantly better than Internet assistants in ”executable ability”. Xiao V shows balanced performance across multiple key dimensions, especially in system control, social interaction, and scene understanding, representing the current mobile terminal large model application level.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Companion of the 2025 ACM International Joint Conference on Pervasive and Ubiquitous Computing; October 12–16, 2025; Espoo, Finlandbooktitle: Companion of the 2025 ACM International Joint Conference on Pervasive and Ubiquitous Computing (UbiComp Companion ’25), October 12–16, 2025, Espoo, Finlanddoi: 10.1145/3714394.3756184isbn: 979-8-4007-1477-1/2025/10doi: 10.1145/3714394.3756184ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Natural language processingccs: Computing methodologies Multi-agent systems
@@ 1. Introduction
  • Cross-platform validation was conducted across five major smartphone brands (vivo, OPPO, Xiaomi, Huawei, and Honor) and three internet-based large-model mobile apps (Doubao, Tencent Yuanbao, and DeepSeek Chat), demonstrating the framework’s real-world applicability and cross-domain generality;
@@ 4.2. Cross-Platform Evaluation Setup
We selected eight representative commercial AI assistant platforms, including five smartphone-embedded systems from leading mobile manufacturers and three internet-based large-model mobile applications, to ensure diversity in model architectures and interaction modalities. For neutrality and confidentiality, these systems are anonymized as Model-A to Model-H in subsequent analyses. To guarantee fairness, all input prompts were delivered through a unified protocol within the same temporal window and device configuration.
@@ 4.4. Results and Analysis
First-order dimension Total Model-A Model-B Model-C Model-D Model-E Model-F Model-G Model-H
@@
Table 2. Each assistant shows distinct performance differences across various scenarios. Models A–C correspond to internet-based large-model applications, while Models D–H represent smartphone-embedded AI assistant systems.
@@
From the perspective of task type dimension (Table 2), each assistant shows obvious differences in different scenarios. In educational learning tasks, the overall performance of each model is better, with average scores generally exceeding 85 points. Internet assistants (such as Model-A 94.82 and Model-C 93.51) show strong multi-modal understanding in image-text problems. Mobile phone assistants demonstrate actual control ability, significantly better than Internet assistants in ”executable ability”. Model-H shows balanced performance across multiple key dimensions, especially in system control, social interaction, and scene understanding, representing the current mobile terminal large model application level.